OpenStudy (freemap):

Which ordered pairs are solutions to the inequality y – 4x ≤ –6? Choose all answers that are correct. A. (–2, –14) B. (–1, –9) C. (0, –5) D. (1, 3) E. (3, 5)

OpenStudy (freemap):

@radar

OpenStudy (freemap):

please help me

OpenStudy (phi):

what is the first (-2,-14) ?

OpenStudy (freemap):

like plugin in -2 for x

OpenStudy (phi):

It is not labeled as a choice? or is it supposed to be choice A ?

OpenStudy (freemap):

yes,

OpenStudy (freemap):

No I mean sorry

OpenStudy (phi):

can you edit the question? otherwise it will be confusing

OpenStudy (freemap):

ok

OpenStudy (freemap):

Which ordered pairs are solutions to the inequality y – 4x ≤ –6? Choose all answers that are correct. A.(–2, –14) B.(–1, –9) C.(0, –5) D.(3, 5) E.(1,3) There all answer choices, pick all that are correct

OpenStudy (phi):

y-4x are "instructions" that says multiply the first number ( the x number in a pair) by -4 then add it to the 2nd number (the y number in a pair) can you do that for the pair (-2,-14) ?

OpenStudy (freemap):

ok, you don't have use answer choices. May i have an example

OpenStudy (phi):

for (-2,-14) multiply the first number by -4: -2*-4 then add the 2nd number: -14 can you do that ?

OpenStudy (freemap):

ok

OpenStudy (freemap):

16--14=30

OpenStudy (phi):

first do -2*-4

OpenStudy (freemap):

8

OpenStudy (phi):

now add -14

OpenStudy (freemap):

-6

OpenStudy (phi):

yes. now we have \[ y - 4x \le -6 \\ -6 \le -6\] the little _ under the \( \lt\) means equal. is -6 = -6 ? yes. so choice A works.

OpenStudy (freemap):

ok, now do the same thing for all of them

OpenStudy (phi):

now try (–1, –9) in -4x+y (this means the same as y + -4x or y-4x) and you do -4* -1 add -9

OpenStudy (freemap):

4+9=13

OpenStudy (phi):

the 2nd number in (-1,-9) is not +9. it is -9

OpenStudy (freemap):

5

OpenStudy (phi):

4+ -9 = what ?

OpenStudy (freemap):

-5

OpenStudy (phi):

when you add numbers with different signs, subtract them (ignoring the sign) and make the answer the same sign as the "bigger" number. yes we get -5

OpenStudy (phi):

now \[ -5 \le -6\] it is clearly not equal. notice the "big" end of the \( \le\) is next to -6. that means we want -6 to be bigger than the other side is -6 bigger than -5 ?

OpenStudy (freemap):

no

OpenStudy (phi):

ok, -5 is bigger than -6 and we want numbers smaller than -6 so cross off choice B. so far , only A works.

OpenStudy (freemap):

ok

OpenStudy (phi):

C. (0, –5) -4 x + y do -4 * 0 then add -5

OpenStudy (freemap):

-5

OpenStudy (phi):

and we just did \( -5 \le -6\). Not true. cross off choice C

OpenStudy (freemap):

ok

OpenStudy (phi):

can you try D. (3, 5) and show your steps.

OpenStudy (freemap):

ok

OpenStudy (freemap):

D. (3,5) -4x + y -4 * 3=-12 add 5 =-7

OpenStudy (phi):

ok, and is \(-7 \le -6\) true or false?

OpenStudy (freemap):

yes

OpenStudy (freemap):

true

OpenStudy (phi):

ok, so choice A and choice D one more to go E. (3, 5)

OpenStudy (freemap):

I thought we just did 3,5

OpenStudy (phi):

wait, E is (1,3)

OpenStudy (freemap):

E. (1,3) -4x+y -4*1=-4 add 3, -4+3=-1 -1 Is bigger than -6 so the equation is false

OpenStudy (phi):

yes. now double check your question, that choice D is (3,5) and E is (1,3) up top you have it the other way round.

OpenStudy (freemap):

It is, Thanks so much, You've been a great help. Could you check another one of my answer. I'll post again

OpenStudy (phi):

ok